# Rearranging small sets for distinct partial sums

### v1

@@ 1. A conjecture of Graham
Although many papers have been written about this conjecture and related problems, the state of the art is still essentially that Conjecture 1.1 holds when $|A|\leq 12$ (see [7] and the references therein) and when $p-3\leq|A|\leq p-1$ (see [12] and the references therein). Many of the arguments for small $A$ use the Polynomial Method and rely on Alon’s Combinatorial Nullstellensatz.
@@ 3. Remarks
    Say that an abelian group $G$ is sequenceable if every subset of $G\setminus\{0\}$ has a valid ordering. The argument of Theorem 1.3 can be easily modified to show the stronger statement that an abelian group $H$ is sequenceable if and only if $H\times\mathbb{Z}$ is sequenceable: In the proof, consider valid orderings of $A$ in which all of the elements with positive second coordinate appear before all of the elements with second coordinate 00, which in turn appear before all of the elements with negative second coordinate. (Theorem 1.3 corresponds to the case where $H$ is the trivial group.) It could be interesting to formulate versions of this principle in nonabelian settings (see [5]).
  4. (4)
    Combining the first and third remarks, we see that if $H_{1}$ is a sequenceable abelian group and $H_{2}$ is an abelian group with no nonzero elements of order strictly smaller than $p$, then every subset $A\subseteq(H_{1}\times H_{2})\setminus\{(0,0)\}$ with $|\pi_{H_{2}}(A)|\leq\log p/\log\log p$ has a valid ordering. In the context of finite cyclic groups, this significantly extends the results of [6] and some of the results of [5].
  5. (5)
@@ Acknowledgements
The author was supported in part by the NSF Graduate Research Fellowship Program under grant DGE–203965. I thank Stefan Steinerberger for bringing Graham’s conjecture to my attention and for engaging in helpful discussions, and I thank Simone Costa for giving useful comments on a draft of this paper.

### v2

@@ 1. A conjecture of Graham
Although many papers have been written about this conjecture and related problems, the state of the art is still essentially that Conjecture 1.1 holds when $|A|\leq 12$ (see [8] and the references therein) and when $A$ is a non-zero sum set of size $p-2$ or $p-3$ (see [13] and the references therein). Many of the arguments for small $A$ use the Polynomial Method and rely on Alon’s Combinatorial Nullstellensatz.
@@
We later learned that Will Sawin [16] proved a very similar result, using the same two main steps, in a MathOverflow post in 2015. His argument and ours differ in the details of both steps, however, and we believe that it is useful to have both approaches recorded in the literature.
@@ 3. Remarks
  2. (2)
    Say that an abelian group $G$ is sequenceable (respectively, strongly sequenceable) if every subset of $G\setminus\{0\}$ has a valid (respectively, two-sided valid) ordering. The argument of Theorem 1.3 can be easily modified to show the stronger statement that an abelian group $H$ is strongly sequenceable if and only if $H\times\mathbb{Z}$ is strongly sequenceable. (Theorem 1.3 corresponds to the case where $H$ is the trivial group.) One can modify the proof of Theorem 1.3 as follows: Partition $A=P\cup Z\cup(-N)$ where $P,N$ are sets of elements with positive second coordinate and $Z$ is a set of elements with second coordinate 00. If $\sum_{z\in Z}z\neq 0$, then consider orderings of $A$ consisting of the elements of $P$, then the elements of $Z$, then the elements of $-N$. If $\sum_{z\in Z}z=0$, then pick some suitable $z^{*}\in Z$ and consider orderings of $A$ consisting of the elements of $P$, then the elements of $Z\setminus\{z^{*}\}$, then the elements of $-N$, with $z^{*}$ placed at either the very beginning or the very end. It could be interesting to formulate versions of this principle in nonabelian settings (see [6]). [^1] Note added in revision: Such results for the infinite dihedral group and other semidirect products with $\mathbb{Z}$ have recently appeared in [5].
  3. (3)
@@
    Combining the second and third remarks, we see that if $H_{1}$ is a strongly sequenceable abelian group and $H_{2}$ is an abelian group with no nonzero elements of order strictly smaller than $p$, then every subset $A\subseteq(H_{1}\times H_{2})\setminus\{(0,0)\}$ with $|\pi_{H_{2}}(A)|\leq\log p/\log\log p$ has a two-sided valid ordering. In the context of finite cyclic groups, this significantly extends the results of [7] and some of the results of [6].
@@ Acknowledgements
The author was supported in part by the NSF Graduate Research Fellowship Program under grant DGE–203965. I thank Stefan Steinerberger for bringing Graham’s conjecture to my attention and for engaging in helpful discussions. I thank Simone Costa for giving useful comments on a draft of this paper, and I thank Raphael Steiner for bringing the reference [16] to my attention.